OpenStudy (anonymous):

Consider the equation v + 4 + v = 8. What is the resulting equation after the first step in the solution? a, v +4 = 8 – v b. v +4 = 8 c. 4 + v = 8 – v d. 2v + 4 = 8

OpenStudy (anonymous):

@Musicdude

OpenStudy (anonymous):

well you have 2 v's what should you do first??

OpenStudy (anonymous):

im not sure @Musicdude

OpenStudy (anonymous):

so if i have x+x+x+4=14 i would change it to 3x+4=14

OpenStudy (anonymous):

you see what i did

OpenStudy (anonymous):

i did what is called "combining like terms" because they both are X i can combine them

OpenStudy (anonymous):

so with those v's what do you think you should do??

OpenStudy (anonymous):

you there?? @destiny5

OpenStudy (anonymous):

i think i should get rid of thee v's @Musicdude

OpenStudy (anonymous):

well in your problem there are two v's

OpenStudy (anonymous):

and we know that v+v=2•v

OpenStudy (anonymous):

so which choice looks like 2v +4

OpenStudy (anonymous):

D? @Musicdude

OpenStudy (anonymous):

Yes :)